OpenStudy (anonymous):

How many possible values for y are there where y =arctan3 ? A. 1 B. 0 C. 2 D. Infinite

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=+y+%3Darctan3+

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

arctan is a function, so arctan(3) is a number

OpenStudy (anonymous):

right? like \[f(x)=x^2+1\] is a function which means that \[f(3)=3^2+1\] is a number. if it was not a single number then \[f(x)=x^2+1\] would not be a function

OpenStudy (anonymous):

maybe the confusion could be that there could be more than one number whose tangent is 3. infinitely many in fact. but arctan(3) is not SOME number whose tangent is 3, it is the number in the interval \[[-\frac{\pi}{2},\frac{\pi}{2}]\] whose tangent is 3